OpenStudy (fanduekisses):

If f(x)= cos^2x, then f''(pi) =

OpenStudy (fanduekisses):

f'(x) = 2*cos(x) * - sin(x), right?

OpenStudy (turingtest):

i would write it as f'(x) = 2*cos(x) * [- sin(x)] to avoid confusion, but yeah

OpenStudy (fanduekisses):

\[f''(x)=(-2\sin(x))(-\sin(x))+(-\cos(x))(2\cos(x))\]

OpenStudy (turingtest):

looking good

OpenStudy (solomonzelman):

yes:)

OpenStudy (fanduekisses):

\[2\sin^2(x) + [-2\cos^2(x)]\]

OpenStudy (turingtest):

so far so good

OpenStudy (solomonzelman):

you can rewirte it is a double angle rule, after factoring out a 2

OpenStudy (solomonzelman):

factor out of \(\large\color{bla k}{ 2 }\), and use: \(\large\color{ k}{ \cos^2(s)-\sin^2(s) =\cos(2s) }\)

OpenStudy (fanduekisses):

so 2(-cos(2x))

OpenStudy (turingtest):

yes, and plug in pi

OpenStudy (fanduekisses):

I get -1.987986 but my options are -2, 0, 1, 2, and 2pi

OpenStudy (fanduekisses):

I guess I'll just round it to -2 :)

OpenStudy (mtalhahassan2):

yep

OpenStudy (turingtest):

cos(2pi) is an exact number

OpenStudy (mtalhahassan2):

Which grade math is it

OpenStudy (solomonzelman):

yes it is:D exact num

OpenStudy (turingtest):



OpenStudy (fanduekisses):

OOh lol the unit circle, yep, maybe I should stay away from calculators lel.

OpenStudy (fanduekisses):

I forgot it was an exact number. D:

OpenStudy (turingtest):

yeah, or if you use them at least remember to make sure you are in radian mode when using pi

OpenStudy (fanduekisses):

Thank you :)

OpenStudy (turingtest):

welcome!

OpenStudy (mtalhahassan2):

which grade math is it

OpenStudy (fanduekisses):

Calculus I

OpenStudy (mtalhahassan2):

ok

OpenStudy (mtalhahassan2):

Is it hard

OpenStudy (mtalhahassan2):

I have same math next semister